Delisile spends 30% if her income on rent and 50% of the remaining on food and travelling. She pays R2 200,00 per month for education and now has R3 120,00 left. Her monthly income is?


1
Expert's answer
2022-02-22T08:41:07-0500

Solution:

Let her monthly income be "R\ x".

Amount spent on rent "=0.3x"

Amount remaining after rent"=0.7x"

Amount spent on food and travelling"=0.5\times0.7x=0.35x"

R 2200 per month for education

Then, "0.35x-2200=3120"

"\Rightarrow 0.35x=2200+3120
\\\Rightarrow 0.35x=5320
\\\Rightarrow x=15200"

So, her monthly income is R 15200.